A certain hydrocarbon on complete combustion at s.t.p produced 89.6dm3 of CO2 and 54g of water. The hydrocarbon should be

A. C6H6
B. C4H10
C. C5H10
D. C4H6

Please show working
1
Expert's answer
2020-04-20T15:08:34-0400

CxHy + O2 = xCO2 + y/2 H2O


If we burn 1 mol of certain hydrocarbon, we would get x mol of CO2 and y/2 mol of H2O.

During combustion, we received 54/18 = 3 mol of water and 89.6/22.4 = 4 mol CO2.

The hydrocarbon should be C4H6 (option D).